OpenStudy (anonymous):

Can you show me how to do this, it makes no sense.(2c/c-4)-2=(4/c+5)

OpenStudy (anonymous):

The answer is -14, but how do you solve that?

hero (hero):

\[\frac{2c}{c - 4} - 2 = \frac{4}{c + 5}\] Right?

OpenStudy (anonymous):

Yes

hero (hero):

Okay \[\frac{2c}{c - 4} - \frac{4}{c + 5} = 2\] Do you know how I got that?

OpenStudy (anonymous):

Yes you added 2 to both sides

hero (hero):

Then I subtracted \(\dfrac{4}{c + 5}\) from both sides.

OpenStudy (anonymous):

Then you subtracted 4/c+5

OpenStudy (anonymous):

Yes

hero (hero):

So now, I have to subtract the fractions, except they don't have the same denominator.

OpenStudy (anonymous):

Correct

hero (hero):

Just a quick question... How would you subtract \[\dfrac{3}{4} - \frac{2}{3}\]?

OpenStudy (anonymous):

Get common denominators then subratct

hero (hero):

Right and how might you go about doing all of that? Can you show your work?